# references.bib

### v2

@@ 5. Discussion
This case study reviewed an effort that combined collective dialogues, bridging-based ranking, and LLMs into a recursive deliberative process that helped find common ground between Israeli and Palestinian peacebuilders. While prior work has demonstrated the effectiveness of integrating similar techniques in low-stakes lab settings, this work demonstrates that these approaches can be useful in high-stakes real-world settings. Critically, the process introduced here was designed to maintain legitimacy in a low-trust environment while still benefiting from using LLMs. To achieve this, the process was designed to accommodate the imperfect outputs from LLMs through a combination of expert humans in the loop, transparency with participants, and final votes to validate (or invalidate) all points of common ground.
@@
LLMs played two critical roles in enabling this process. First, they enabled rapid distillation of many raw statements into a manageable set of collective statements that captured the main ideas while maintaining their linguistic nuances. This supplanted the complex, multi-stakeholder deliberation common in politically sensitive synthesis that can take months, with a LLM-accelerated step that took a few hours. As a result, each dialogue cycle could be compressed to a few weeks, making the timeline for a recursive process with many cycles viable. Second, LLMs and machine translation made the multilingual cycle practical. LLMs decreased the burden of complex translation tasks and enabled multi-lingual collective statement synthesis, while low latency machine translation made the live multi-lingual collective dialogues viable.
5.1. Ethical Considerations
@@
Undertaking an exercise such as this in the context of active conflict raises many valid ethical considerations. We summarize some of those we considered below, along with the steps taken to address them. In general, when making decisions about delaying or potentially not running the process, we weighed both the potential downside and upside of moving forward against the “null hypothesis” of the status quo.
@@
  • Backfire: There was a risk that conducting the large-scale cross-border dialogue in the joint phase could backfire in various ways (e.g., if dialogue broke down), and inflame the division this effort sought to bridge. We mitigated this risk by starting with a uninational phase designed to help build trust and set the conditions for a productive cross-border dialogue.
@@
    • The use of imperfect AI and ML algorithms throughout the process risks injecting errors or biases that misrepresent participants. We mitigated this risk via humans in the loop: experts reviewed and refined all outputs downstream of AI/ML, and final votes among participants were conducted to validate all potential points of common ground. To minimize bias from errors in machine translation or LLM processing, we made LLM pipeline choices using micro experiments that enabled native Hebrew and Palestinian Arabic speakers decide what would be used for their language.
@@
5.2. Limitations
On the one hand, this use case might be viewed as a success, as it demonstrated an approach that indeed produced points of common ground amid tense real-world division. However, several limitations deserve mention:
@@ 5.2. Limitations
  • Participant selection bias: The study participants were self-selected peacebuilders and thus are likely not be representative of the broader Israeli and Palestinian populations. This limits the generalizability of the findings to the wider public, who may hold more entrenched or divergent viewpoints.
@@
  • Technological dependency: The reliance on online collective dialogue and LLMs tools may introduce a selection bias based on technological literacy and access. The success of this type of process is contingent on reliable internet access, which is not universally available in all settings or to all groups.
5.3. Future Directions
@@
  • Going beyond distillation: AI can be used not just for distilling information, but for generating novel more exploratory proposals. Using an iterative deliberation method, similar to the one proposed by Tessler et. al (tessler2024ai), an LLM could generate a set of initial collective statements that people then critique and refine. This process could explore novel ideas that emerge through a process of iterative synthesis and refinement, with a focus on balanced perspectives. These combined ideas might not individually represent common ground but, when combined in a balanced way, could reflect a form of reasoned consensus. Such statements could not be derived solely from individual statements with high bridging metrics. It would be essential that the resulting statements from this kind of generative process still be validated with a final vote of agreement among human participants to ensure that they actually represent common ground.
@@
Acknowledgements.
We first thank the team at ALLMEP that worked extremely hard to make this effort happen: Doubi Schwartz, Wasim Almasri, Nivine Sandouka, Brian Reeves, Nick Acosta, Kari Reid, Dr. Natali Levin-Schwartz, John Lyndon, and Avi Meyerstein; Avi and John for their leadership and commitment to advancing this initiative during challenging times. We thank Hope Schroeder, Manuel Wuthrich, Karmel Salah, Jay Baxter, Colin McGill, and Colin Irwin for helpful conversations and input. Finally, we thank Jake Weiner, Ivy Cheung, Amanda Yonce, Margo Flewelling, Jennifer Gong, and Emmet Hennessy for rapidly developing the multi-lingual capability on Remesh that enabled timely execution of the joint phase. Luke Thorburn was supported in part by UK Research and Innovation [grant number EP/S023356/1], in the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence (safeandtrustedai.org), King’s College London.
[…]
Appendix A Prompts
A.1. Used for dialogue cycle 1 in Hebrew
A.2. Used for dialogue cycles 2 and 3 in Palestinian Arabic
A.3. Used for dialogue cycle 4 in Palestinian Arabic, Hebrew, and English

### v3

@@ 5. Discussion
In this section, we discuss how this work can be evaluated, and considerations and challenges encountered when using AI in this context. Further discussion of ethical issues, limitations, and future directions are included in Appendix D.
Evaluation
As this project was conducted in support of an active peace process, it lacked the structure of well-controlled trials that support rigorous comparative evaluations. Further, our process did not directly replace an existing one to which it could be directly compared. To the extent that it was replacing anything, it could be partly viewed as replacing long deliberations where professional mediators employ standard, labour-intensive approaches to finding common ground, e.g., (Moore, 2014; Dillard, 2013; Schirch and Campt, 2015). In contrast, ALLMEP noted that “Without this process, our conversations would be smaller and fragmented. This method brought voices together across political and geographic divides in a way that traditional dialogue simply couldn’t.
The 10 public-oriented statements produced by this process had between 84-96% bridging support, and nine had support over $90\%$. While no perfect analogue exists to say how “good” these levels of support are, we note that citizens assemblies — often cited for their ability to find common ground on divisive issues — generally produce recommendations that have 60-95% overall support under less tense circumstances (British Columbia Citizens’ Assembly on Electoral Reform, 2004; Farrell et al., 2015; BürgerBegehren Klimaschutz, 2020; Gabbatiss, 2020; Global Assembly, 2022; Involve, 2018).
Participants in the process reported that the experience fostered ”a lot of self-reflection,” that the anonymity enabled them to ”share openly, without fear of being judged” and that ”seeing each other’s anonymous thoughts in this way is valuable in and of itself.” However, while most participants reported a positive experience, some thought that in the live collective dialogues ”Everything is so fast. It’s hard to think.” and that they ”need to be able to speak to each other, see each other’s faces” (see Appendix E for more of this feedback).
Use of AI
@@
LLMs played a few critical roles in enabling this process. First, they rapidly distilled numerous raw statements into manageable, nuanced collective statements, replacing complex multi-stakeholder deliberations (which can take months) with an LLM-accelerated step of a few hours. This compressed each dialogue cycle to weeks, making a recursive process viable. Second, LLMs and machine translation made the multilingual cycle practical by reducing complex translation burdens, enabling multilingual collective statement synthesis, and facilitating live multi-lingual dialogues through low-latency translation. Third, LLM-enabled vote prediction (Konya et al., 2022) improved data efficiency for the bridging-based ranking and decreasing how much participants had to vote per prompt and thus allowing more prompts within time-constrained dialogues.
A key challenge of this work was figuring out how to reap these benefits of using AI without sacrificing trust and legitimacy due to real/perceived issues with AI. Vote prediction is imperfect, LLMs can hallucinate, and machine translation is sometimes just wrong. Such imperfections hurt trust and legitimacy in a system that uses AI. However, rather than ensuring all AI used was unbiased, and error-free, our approach was designed under the assumption that all AI used would be imperfect. The process was formulated to accommodate AI imperfection by incorporating expert humans in the loop, being transparent with participants (Appendix C.1), and using AI-free final votes to validate all points of common ground. In particular, all outputs of AI/LLMs were verified with humans-in-the-loop, with the exception of real-time machine translation during the dialogues for which we had the experts review sample translations in advance to choose the most accurate translation method available. AI was only used for tasks where it was faster or cheaper to have the AI do it and then have humans check the outputs, than to have the humans do it themselves.
6. Conclusion
In this case study, we documented a recursive deliberative process in which we used online collective dialogues, bridging-based ranking, and LLMs to surface common ground between Israeli and Palestinian peacebuilders. While prior work has shown the effectiveness of integrating similar techniques in low-stakes settings, this work is an “existence proof” that (i) these technologies can be used within high-stakes conflict settings while maintaining trust; (ii) that a scalable method can successfully bridge divides among distrustful groups amid active conflict; and (iii) that such processes can be made equitable despite significant structural asymmetries between participants.
Acknowledgements.
We first thank the team at ALLMEP that worked extremely hard to make this effort happen: Doubi Schwartz, Wasim Almasri, Nivine Sandouka, Brian Reeves, Nick Acosta, Kari Reid, Dr. Natali Levin-Schwartz, John Lyndon, and Avi Meyerstein; Avi and John for their leadership and commitment to advancing this initiative during challenging times. We thank Hope Schroeder, Manuel Wuthrich, Karmel Salah, Jay Baxter, Colin Megill, and Colin Irwin for helpful conversations and input. Finally, we thank Jake Weiner, Ivy Cheung, Amanda Yonce, Margo Flewelling, Jennifer Gong, and Emmet Hennessy for rapidly developing the multi-lingual capability on Remesh that enabled timely execution of the joint phase. Luke Thorburn was supported in part by UK Research and Innovation [grant number EP/S023356/1], in the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence (safeandtrustedai.org), King’s College London.
Appendix A Joint Letter
“We write as a united body of Israeli and Palestinian peacebuilders, compelled by the urgency of our shared pain and galvanized with the hope we still hold for the future. We call for the following:
  • An immediate ceasefire and release of all hostages
  • The creation of an international fund to build the foundations for a comprehensive long-term solution
  • The establishment of a globalized coalition to assist Gaza’s reconstruction
  • The inclusion of civil society and NGOs in all peacebuilding and diplomatic discussions
  • The recognition of both sides’ trauma
Acknowledging the trauma and pain experienced by both sides is essential for the healing of our societies and the achievement of sustainable peace. Help us break the cycle of violence so the next generation may flourish. The world watches, and history will judge. Now is the time to act.”
Appendix B Further Background
B.1. Collective Dialogue Dynamics
During a live collective dialogue process, all participation is simultaneous and each collective response process takes a few minutes. When each response process completes, each participant sees the fraction of all participants agreeing with their response along with a representative subset of responses from the group. The moderator sees live results which they can use to decide if they need to pivot from the pre-programmed discussion guide to ask additional questions on an issue. After a live dialogue is complete it can be left open for asynchronous participation so those who could not make the live event can still participate. Collective dialogues can also be run asynchronously from the start if the logistics of a live dialogue are not feasible for the situation.
Appendix C Further Methodological Details
C.1. Transparency about the use of AI
In the invitations and public communications, participants were told “AI” was being used, e.g., “we are using an AI-based tool to hold a series of large scale ‘collective dialogues’ among Israelis and Palestinians.” F Some participants attended optional office hour/Q&A calls, in which we discussed how LLMs specifically were being used.
Most prospective participants responded positively. Some were skeptical of the use of AI and wanted to understand the details of how it was being used further. For those, we held a few ‘office hour’ sessions where we walked anyone interested or concerned through the details of how we used AI. Most questions were around elicitation inference & how well LLMs could handle Hebrew/Arabic. We typically walked them through the overall process, explained the micro-experiments for the LLM pipeline related to Hebrew/Arabic, and shared links to the elicitation inference papers.
It’s unclear exactly how awareness of the use of LLMs shaped the consensus-building process, but we think that talking to a “computer” made it easier for people to express themselves freely. For some, LLMs were seen as more impartial. For others, there was persistent skepticism.
Appendix D Further Discussion
D.1. Ethical Considerations
@@
Undertaking this work this in the context of active conflict raises many valid ethical concerns. We summarize some of those we considered below, along with steps taken to address them. In general, when making decisions about delaying or potentially not running the process, we weighed the potential downsides and upsides of moving forward against the “null hypothesis” of the status quo.
@@
  • Backfire: There was a risk that conducting the cross-border dialogue in the joint phase could backfire in various ways (e.g., if dialogue broke down), and inflame the division this effort sought to bridge. We mitigated this risk by starting with a uninational phase designed to help build trust and set the conditions for a productive cross-border dialogue.
@@
    • The use of imperfect AI and ML algorithms throughout the process risks injecting errors or biases that misrepresent participants. We mitigated this risk via humans in the loop: experts reviewed and refined all outputs downstream of AI/ML, and final votes among participants were conducted to validate all points of common ground. To minimize bias from errors in machine translation or LLM processing, we made LLM pipeline choices using micro experiments that enabled native Hebrew and Palestinian Arabic speakers decide what would be used for their language.
@@
D.2. Limitations
Several limitations of this case study deserve mention:
@@ D.2. Limitations
  • Participant selection bias: The study participants were self-selected peacebuilders and thus not be representative of the broader Israeli and Palestinian populations. It remains unclear if the approach will generalize of to a wider public who may hold more entrenched or divergent viewpoints.
@@
  • Technological dependency: The reliance on online collective dialogue and LLMs tools may introduce a selection bias based on technological literacy and access. The success of this type of process is contingent on reliable internet access, which is not (yet) universally available.
D.3. Future Directions
@@
  • Going beyond distillation: AI can be used not just for distilling information, but for generating novel more exploratory proposals. Using an iterative deliberation method, similar to the one proposed by Tessler et. al (Tessler et al., 2024), an LLM could generate a set of initial collective statements that people then critique and refine. This process could explore novel ideas that emerge through a process of iterative synthesis and refinement, with a focus on balanced perspectives. These combined ideas might not individually represent common ground but, when combined in a balanced way, could reflect a form of reasoned consensus. Such statements could not be derived solely from individual statements with high bridging metrics. It would be essential that the resulting statements from this kind of generative process still be validated with a final vote of agreement among human participants to ensure that they actually represent common ground.
@@
Appendix E Qualitative Feedback
E.1. ALLMEP Retrospective
Below are a few specific questions asked to ALLMEP leaders involved in the process. Direct quotes denote their verbatim responses and content not in quotes is context provided by this paper’s authors.
  • Did you find [the process] helpful?
    “Yes. At a time when real dialogue felt impossible post October 7th, this process allowed Palestinians and Israelis to engage without compromising their experiences or red lines.”
  • Do you plan to use the process again?
    “Absolutely. The ability to engage anonymously and equitably is critical, especially for Palestinians who face political and social risks when participating in cross-border work.”
    (We are continuing to work with ALLMEP on scaling this process up to larger and more representative samples of the public; of which — as of writing this — the first dialogue cycle has already begun.)
  • Were there parts of the process that you were especially excited about or skeptical of [before it was run]?
    “I was mostly excited about the uninational phase. It gave [us] room to define our priorities before entering joint dialogue, which felt empowering and necessary. I was a little skeptical however whether AI could reflect our political realities.”
    Beyond all the trust building aspects discussed throughout the paper, we note that using the well-known Community Notes algorithm helped give the process improved face validity and build trust.
E.2. Participant Feedback
E.2.1. Examples of Positive Feedback
  • “I found it to be very validating and very positive! It gave me hope”
  • “That was an INCREDIBLE experience”
  • “I felt that I could share openly, without fear of being judged. It felt much less complicated than sharing directly with colleagues”
  • “This was an interesting experience, a lot of self-reflection regarding our inner motives for this work”
  • “I know there is an end goal for this, but seeing each other’s anonymous thoughts in this way is valuable in and of itself and perhaps this sort of work could be integrated in various ways as part of more regular activities. The openness of anonymity is important.”
E.2.2. Examples of Negative or Ambivalent Feedback
  • “Interesting activity but time for answers was a bit short”
  • “Ummm…. I haven’t made up my mind about this technology yet….”
  • “Everything is so fast. It’s hard to think. It was possible to give a little more time to read and respond. The idea is cool”
  • “We need to be able to speak to each other, see each other’s faces”
  • “I am not sure that I was able to express myself completely or clearly also due to technical limitations of time and the translation.”
Appendix F Recruitment Invitation
Appendix G Sample Sizes
Table 4. The tables below record what we know about the number of participants in each dialogue cycle. Columns “CDa”, “CDb”, “CDc” respectively contain (a) the number who joined the collective dialogue and did demographic onboarding, (b) the number who simultaneously contributed a statement to the same open-ended prompt, and (c) the number contributed statements and voted all the way through to the last collective prompt. Columns “Va”, “Vb”, “Vc” respectively contain (a) the number who joined and did demographic onboarding for the asyncronous vote, (b) the number who additionally voted on all collective statements in the paper (joint) or on 50% of collective statements (uninational), and (c) voted on all collective statements not shared in the paper related to internal peacebuilding strategy. Finally, column “MUP” contains the maximum possible number unique participants of any sort during the dialogue cycle (Va + CDa), and “UEP” contains the best estimate of unique engaged participants during the dialogue cycle (max(CDb, Vb)).
Appendix H Prompts
H.1. Used for Dialogue Cycle 1 in Hebrew
H.2. Used for Dialogue Cycles 2 and 3 in Palestinian Arabic
H.3. Used for Dialogue Cycle 4 in Palestinian Arabic, Hebrew, and English